If mass of a neutron is double ,Mass of electron is double .Find out the new atomic weight of 6C12
 

Grade:12

1 Answers

Vikas Amritiya
2 years ago
6​C12→e=6
Number of protons =6=6amu
Number of neutrons =6=6amu
Total mass =(6+6)=12amu
Mass of the electron is negligible.
If the mass of the neutron is doubled, then the total mass =(6×2)+6=18amu
Therefore, the % of the increase ==1218−12​×100=50%